Out of 500 people sampled, 235 preferred Candidate A. Based on this, find a 95% confidence level for the true proportion of the voting population (P) prefers Candidate A. Give your answers as decimals, to three places <pく 20 Points possible:1
Answer #1

Solution :

Given that,

n = 500

x = 235

\hat p = x / n = 235 / 500 = 0.470

1 - \hat p = 1 - 0.470 = 0.530

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 1.96 * (\sqrt((0.470 * 0.530) / 500)

= 0.044

A 95% confidence interval for population proportion p is ,

\hat p - E < P < \hat p + E

0.470 - 0.044 < p < 0.470 + 0.044

0.426 < p < 0.514